Determine the momentum of a (4.600x10^5) kg airplane traveling southwest at (6.1x10^2) km/h.

Note: Your answer is assumed to be reduced to the highest power possible.

Your Answer:


1
Expert's answer
2021-10-25T15:24:15-0400
"p=mv,""p=4.6\times10^5\ kg\times6.1\times10^2\ \dfrac{km}{h}\times\dfrac{1000\ m}{1\ km}\times\dfrac{1\ h}{3600\ s},""p=7.8\times10^7\ \dfrac{kg\times m}{s}."